# Machine learning Hubbard parameters with equivariant neural networks

### v1

@@ RESULTS
In the following, for the sake of convenience, we label these tensors in their irrep form as $\mathbf{x}^{i}_{\mathrm{p}}$ for $p$ orbitals and $\mathbf{x}^{i}_{\mathrm{d}}$ for $d$ orbitals, where $i$ can be 1 or 2. We note in passing that in the non-spin-polarized case, one can set $\mathbf{n}_{\ell}^{\uparrow}=\mathbf{n}_{\ell}^{\downarrow}=\mathbf{n}_{\ell}/2$, where $\mathbf{n}_{\ell}$ is divided by two to account for spin degeneracy, and then proceed as above. In addition, atomic species (i.e. atomic types) are encoded as one-hot vectors in our ML model. For example, for a model that supports $N_{s}$ distinct atomic species, we choose an arbitrary labelling of a $N_{s}$ long vector, $[\text{Mn},\text{Fe},\text{Ni},\text{O},\ldots]$, and place a 1 at the entry corresponding to the site element and zeros everywhere else (Mn = $[1,0,0,0,0,\ldots]$, Fe = $[0,1,0,0,0,\ldots]$, etc). A summary of the irreps of all the model inputs is presented in Table 1. In the case of the ML model for Hubbard $V$, the interatomic distances $r_{JI}$ are also included as an additional input, giving some basic information of the ionic structure.
@@
To construct our ML model, we use the e3nn library [73, 75] and PyTorch; [76] our codebase is open-source and freely available (see code availability below). As shown in fig. 2, we define two separate models, one for predicting the on-site $U$ values and the other for the inter-site $V$ ones. By using separate models, we provide flexibility for the model to be used in calculations where only Hubbard $U$ is applied (i.e. DFT+$U$). In each case, the model starts with one or more nodes that carry attributes, expressed as a direct sum of irreps, which represent the inputs to the learned function. The inputs pass through a series of repeated layers made up of a tensor product (analogous to all-to-all connected layers in a traditional neural network) and a gated non-linearity (an equivariant version of a traditional activation function [75]) followed by a final tensor product before readout. All of the tensor products have learnable weights, meaning that every pair of input irreps that contribute to one output irrep has a learnable scalar parameter that is optimised during training. We use the AdamW [77] optimiser to minimise the loss, which is the mean squared error between predicted and training Hubbard parameters.
@@
To train and validate our ENN model, we curated a dataset comprising materials with diverse crystal structures. This dataset is constructed based on an investigation of various Li-ion cathode materials, covering olivine-type, [18] spinel-type, [52] and layered-type [78] structures at different Li concentrations, employing the self-consistent DFT+$U$ and DFT+$U$+$V$ protocols outlined above. Additionally, we include materials such as tunnel- and rutile-type MnO2 [79, 51] and perovskite-type rare-earth nickelates [53] in our dataset. The crystal and electronic structure details, magnetic ordering, and various other properties of these materials are extensively documented in their respective publications. These are the compounds containing TM elements Ni, Mn, and Fe, and a list of all these materials is provided in table 2. It is crucial to note that, while our current dataset is relatively small in terms of material count, ongoing efforts are directed towards a high-throughput exploration involving hundreds of materials, which will serve as an expanded training and validation set for our ENN model in the future. As a proof of concept, this initial dataset demonstrates the efficacy of our ML model, showcasing its ability to predict Hubbard parameters even with a limited dataset. Despite the size of the dataset, the model exhibits accurate predictions, setting the stage for further refinement and expansion with the upcoming comprehensive investigation.
@@
Total 27
Table 2: A list of the materials that are used to train and validate the ML model. The last column shows the number of materials. For the olivine-type materials $x=0,0.25,0.50,0.75,0$, while for the spinel- and layered-type materials $x=0,1$.
For each material in our dataset, SC Hubbard $U$ and $V$ parameters are computed using the self-consistent protocol illustrated in fig. 1(a). We include all $V$ parameters for Hubbard-active atom pairs whose DFPT values are greater than $0.3$ eV, while in Ref. 62 only $V$ for the nearest neighbor couples were included in the training of their ML model. Throughout the iterative process, all intermediate converged occupation matrices and Hubbard parameters are systematically saved. This dataset is then used for training our ENN model. The training data, as summarized in table 3, is quite substantial when considering the multitude of these first-principles calculations. The dataset for training $V$ comprises approximately seven times more data points than that for training $U$ due to the greater abundance of atom pairs involved in the computation of $V$, in contrast to the local and consequently sparser nature of $U$. For each material and chemical element, this data is randomly split, holding back 20% from each material for validation.
@@
Table 3: Chemical elements and the total number of data points for both on-site $U$ and inter-site $V$ Hubbard parameters determined using the SC protocol illustrated in fig. 1(a). These data points are aggregated across all the materials listed in table 2 that contain the respective chemical element. Numbers in brackets indicate the unique data points after de-duplication (see text).
In the first-principles calculations, certain input Hubbard parameters ($U_{\mathrm{in}}$ and $V_{\mathrm{in}}$) yield a specific ground state and occupation matrices, and the response of the system due to a perturbation through DFPT provides corresponding output Hubbard parameters ($U_{\mathrm{out}}$ and $V_{\mathrm{out}}$). This one-to-one correspondence allows our ML model to discern the mapping between input and output Hubbard parameters and the associated intermediate occupation matrices. In addition, it is worth noting that many of the structures encompass symmetry-equivalent sites, and correlations between results from successive self-consistent steps are prevalent. To address this issue, we implement a de-duplication procedure aimed at mitigating potential biases in our training data. This involves calculating a symmetry-invariant distance between all pairs of ENN inputs for each atomic species in each material, which is then used to cluster duplicates based on a specified distance threshold (see Sec. S4 in the SI). In random train/validate splits, we sample from these clusters rather than individual data points, ensuring that only one example from each cluster is included in the validation set.
@@
First, we aim to learn the final SC Hubbard parameters for the TM elements across all the materials in our dataset. This is achieved using all attributes listed in table 1 as inputs for our ML model. Figure 3 shows the parity plots obtained for the on-site $U$ and inter-site $V$ Hubbard parameters. The mean absolute relative error (MARE) over all species consistently remains below 3% and 5% for the $U$ and $V$ parameters, respectively (per-species distributions are reported in Sec. S5 in the SI). The result for Hubbard $U$ is particularly promising, considering the relatively small amount of available training data. The higher MARE obtained for $V$ is likely due to the significantly increased number of degrees of freedom in its model, owing to the involvement of pairs of atoms. It is worth noting that in order to attain this accuracy in predicting the $U$ and $V$ values, utilizing a relatively small batch size of $8-16$ was beneficial in preventing overfitting.
Figure 3: Parity plots showing the prediction accuracy on an unseen validation dataset, where the energies in the legend are the RMSE categorized by element(s) and the overall RMSE across all elements. All attributes listed in table 1 are used as inputs for the ML model. (a) Hubbard $U$ for 3d states of TM elements, (b) Hubbard $V$ between the TM-3d and O-2p states.
Second, we examine the extent to which our ML model depends upon the input Hubbard parameters for its prediction accuracy. To investigate this, we repeat the same numerical experiment as above, eliminating the input Hubbard parameters ($U_{\text{in}}$ and $V_{\text{in}}$) from our model. As detailed in Sec. S6 in the SI, the parity plots closely resemble those in fig. 3. The reduction in accuracy of our ML model resulting from this simplification is relatively small, namely the overall root-mean-square error (RMSE) is increases by 29 and 18 meV for the on-site $U$ and inter-site $V$, respectively. This finding suggests that the input Hubbard parameters may not be the most critical input attributes for our ML model, and that the input occupation matrices contain sufficient information to make low-error predictions.
@@
Table 4: Comparison of the open-circuit voltages $\Phi$ (in V) and magnetic moments for TM elements (in $\mu_{\mathrm{B}}$) for LixMnPO4 and LixFePO4 computed within DFT+$U$+$V$, with $U$ and $V$ obtained from first principles using DFPT and predicted using the ML model. The voltages are computed using the total energy differences for the Li concentrations $x=0$ and $x=1$, while the magnetic moments are computed as the trace of the difference between the spin-up and spin-down occupation matrices. [18] $\Delta\Phi$ and $\Delta m$ are the relative differences between the voltages and magnetic moments based on DFPT and ML, respectively.
Although the model’s relative error is low, it is important to evaluate its impact on downstream properties, such as e.g. voltages in Li-ion battery cathode materials. Table 4 shows a comparison of calculated open-circuit voltage (OCV) [18] within the DFT+$U$+$V$ framework, using first-principles SC Hubbard parameters from DFPT and those predicted by our ML model. We find that the differences between the computed and predicted $U$ and $V$ values are less than $0.12$ eV for LixMnPO4 and LixFePO4 ($x=0$ and $x=1$). Generally, such a small variation in the values of Hubbard parameters has a negligible impact on the vast majority of various physical and chemical properties of materials. The differences observed in the OCV in Table 4 are indeed negligible, indicating the accuracy and reliability of the ML-predicted Hubbard parameters.
@@
In the preceding section, we assessed the ML model’s performance in predicting final self-consistent Hubbard parameters using training data that consists of a subset of DFPT calculations from all iterations of the SC protocol (see fig. 1(a)). However, this approach relies on conducting numerous computationally intensive DFPT calculations to generate these training data (often $2-5$ but occasionally up to $10$ per material to reach self-consistency). In this experiment, we assess the performance of the model by training on only the first $N_{\mathrm{iter}}-1$ DFPT result, and asking it to predict the outcome of the next DFPT calculation. This gives an indication of whether the model can be useful in a scenario where a limited number of self-consistent steps are carried out, such as during the early stages of a high-throughput screening study.
@@
Figure 4 shows the average RMSE over all elements as the number of SC iterations used for training $N_{\mathrm{iter}}$ increases. The reference data corresponds to RMSE evaluated using DFPT calculations, i.e. excluding the ML component. This RMSE is computed by determining the difference between the Hubbard parameters at the current $N_{\mathrm{iter}}$th and the previous $(N_{\mathrm{iter}}-1)$th iteration. It is evident that the reference RMSE decreases non-monotonically, which was observed in previous studies, [19] and eventually diminishes at large $N_{\mathrm{iter}}$. It is noteworthy that the optimal $N_{\mathrm{iter}}$ for achieving self-consistency in the SC protocol varies for different materials listed in table 2. Therefore, for larger $N_{\mathrm{iter}}$ values in fig. 4, fewer data points are available compared to smaller $N_{\mathrm{iter}}$ values.
Figure 4: RMSE as a function of the number of iterations $N_{\mathrm{iter}}$ in the SC protocol (see fig. 1(a)). The reference data represents the RMSE obtained from successive DFPT calculations (without ML involvement), whereas the ML data denotes the RMSE resulting from ML predictions for the $N_{\mathrm{iter}}$th iteration based on training the model on all preceding $N_{\mathrm{iter}}-1$ iterations, with validation conducted using the DFPT data for the $N_{\mathrm{iter}}$th iteration as a reference.
@@
The ML-based RMSE in fig. 4 at the $N_{\mathrm{iter}}$th iteration is computed by training the model on data from all previous $N_{\mathrm{iter}}-1$ iterations and predicting the Hubbard parameters for the $N_{\mathrm{iter}}$th iteration. Section S7 in the SI contains parity plots for various values of $N_{\mathrm{iter}}$, ranging from 2 to 7, alongside the reference results broken down by element as well as the results for the training set. It is evident from fig. 4 that the ML-based RMSE steadily decreases, and for all $N_{\mathrm{iter}}$ values except $N_{\mathrm{iter}}=6$, it is smaller than the RMSE of the reference dataset. This result is shows that our ML model can improve upon a DFPT-based SC result that was terminated early, particularly during the first few iterations, thereby facilitating faster convergence of Hubbard parameters. For instance, the reference value for $N_{\mathrm{iter}}=2$ indicates that on average, a single-shot DFPT calculation yields Hubbard parameters with an RMSE of $\sim 1$ eV, while the ML model trained on single-shot data can reduce this error to under $0.8$ eV. Incorporating additional $N_{\mathrm{iter}}$ iterations in the ML training data leads to steady improvements, although the relative performance compared to the reference gradually diminishes until reaching the floor of the model’s accuracy. This floor is partly determined by the convergence of the underlying DFPT-based Hubbard parameters in the SC protocol, conducted with a tolerance of approximately $\Delta=0.01-0.1$ eV (see fig. 1(a)). Therefore, the most significant computational savings in predicting the SC Hubbard parameters can be achieved using our ML model, which acts as a surrogate, after only a few iterations in the DFPT-based SC protocol. This substantially reduces the overall computational cost and renders it a feasible approach, particularly for high-throughput screening scenarios where a priori training data is unavailable for all materials.
@@
The outcomes of our numerical experiments are shown in fig. 5. As a reference, we compute the RMSE individually for each material by training the ML model on 80% of the data from all five concentrations and validating it on the remaining 20% of the data after de-duplication. This yields RMSE values of 126, 182, and 420 meV for the Fe, Mn, and mixed Fe-Mn olivines, respectively (see the horizontal lines in fig. 5). Subsequently, we investigate the RMSE values for scenarios in which the ML model is trained on fewer concentrations, denoted as $N_{\mathrm{c}}$, to assess the sensitivity of the ML model to the amount of the training data and its transferability for predicting Hubbard parameters at other concentrations. To accomplish this, we train the ML model on $N_{\mathrm{c}}$ concentrations and validate it on the remaining $5-N_{\mathrm{c}}$ concentrations, and then we average out over all possible permutations of the concentrations between the training and validation datasets. The average RMSE values are represented by dots, while the error bars in fig. 5 indicate the maximum and minimum RMSE values resulting from the various permutations. It can be seen that as $N_{\mathrm{c}}$ increases, the average RMSE for each material decreases, reaching values of approximately 106, 301, and 652 meV at $N_{\mathrm{c}}=4$ for the Fe, Mn, and mixed Fe-Mn olivines, respectively.
This experiment demonstrates that, depending on the desired accuracy, it is not necessary to include training data from the exact compositions that the model is being evaluated on, and even using examples at one or two concentrations may be sufficient to extrapolate to other parts of the composition range. Given the computational cost of SC calculations, this can lead to a significant speedup, particularly in cases requiring larger supercells and various configurations of Li ion distributions within the structure. Overall, these findings demonstrate the good transferability of our ML model, indicating its potential utility and reliability for predicting Hubbard parameters in materials not included in the model’s training dataset.
(a) LixFePO4 (Ref. $=0.126$ eV)
(b) LixMnPO4 (Ref. $=0.182$ eV)
(c) LixFe0.5Mn0.5PO4 (Ref. $=0.420$ eV)
@@
Figure 5: RMSE as a function of $N_{\mathrm{c}}$ (the number of Li concentrations $x$) for the three olivines: (a) LixFePO4, (b) LixMnPO4, and (c) LixFe0.5Mn0.5PO4. The ML model is trained and validated using $N_{\mathrm{c}}$ and the remaining $5-N_{\mathrm{c}}$ concentrations, respectively. Error bars indicate the range of RMSE values obtained by considering various permutations of concentrations in the training and validation datasets, the dots represent the average. Reference data is the RMSE computed by training the ML model on 80% of all five concentrations, with validation performed on the remaining 20% after de-duplication. The line represents the mean results over three runs with different random initializations, the confidence interval show the standard deviation.

### v2

@@ RESULTS
In the following, for the sake of convenience, we label these tensors in their irrep form as $\mathbf{x}^{i}_{\mathrm{p}}$ for p orbitals and $\mathbf{x}^{i}_{\mathrm{d}}$ for d orbitals, where $i$ can be 1 or 2. We note in passing that in the non-spin-polarized case, one can set $\mathbf{n}_{\ell}^{\uparrow}=\mathbf{n}_{\ell}^{\downarrow}=\mathbf{n}_{\ell}/2$, where $\mathbf{n}_{\ell}$ is divided by two to account for spin degeneracy, and then proceed as above. In addition, atomic species (i.e. atomic types) are encoded as one-hot vectors in our ML model. A summary of the irreps of all the model inputs is presented in table 1.
In the case of the ML model for Hubbard $V$, the interatomic distances $r_{IJ}$ are also included as an additional input, giving some basic information on the ionic structure. It is important to note that we exclusively focus on the on-site occupation matrices $\mathbf{n}_{\ell}^{II\sigma}$ for practical reasons. While inter-site matrices (see e.g. 78, 79) could be naturally included in our models, Quantum ESPRESSO does not currently output these. Additionally, this approach ensures consistency with the linear-response approach used to calculate the Hubbard $V$ parameters from first principles, which relies on the response of on-site occupation matrices only 9, 48. However, using inter-site occupation matrices $\mathbf{n}_{\ell}^{IJ\sigma}$ as descriptors instead of $r_{IJ}$ presents an interesting alternative that could be explored in future studies. Finally, while additional descriptors, such as the derivatives of the occupation matrix with respect to the Hubbard parameters, could potentially enhance the performance of the ML model, this would significantly increase the cost of generating training data.
Figure 2: Schematic illustration of the equivariant neural-network ML model for predicting $U$ and $V$ Hubbard parameters. $U_{\mathrm{in}}$ and $V_{\mathrm{in}}$ are the input Hubbard parameters, while $U_{\mathrm{out}}$ and $V_{\mathrm{out}}$ are the outputs. The atomic species enter as one-hot tensors, and $r_{IJ}$ is the interatomic distance between sites $I$ and $J$. $\mathbf{x}^{1}_{\mathrm{d}}$ and $\mathbf{x}^{2}_{\mathrm{d}}$ are the occupation matrices for the d orbitals, while $\mathbf{x}^{1}_{\mathrm{p}}$ and $\mathbf{x}^{2}_{\mathrm{p}}$ for the p orbitals, all in the special representation for the ML model [see eqs. 5 and 6]. Tensors are represented as squares: open for inputs and outputs, and filled for intermediate features.
@@
To construct our ML model, we use the e3nn library 75, 77 and PyTorch; 80 our codebase is open-source and freely available (see code availability below). As shown in fig. 2, we define two separate models, one for predicting the on-site $U$ values and the other for the inter-site $V$ ones. By using separate models, we provide flexibility for the model to be used in calculations where only Hubbard $U$ is applied (i.e. DFT+$U$). In each case, the model starts with one or more nodes that carry attributes, expressed as a direct sum of irreps, which represent the inputs to the learned function. The inputs pass through a series of repeated layers made up of a tensor product (analogous to all-to-all connected layers in a traditional neural network) and a gated non-linearity (an equivariant version of a traditional activation function 77) followed by a final tensor product before readout. For each experiment, we rescale the model outputs to have zero mean and a standard deviation of one based on the Hubbard parameters found in the training set which simultaneously accelerates training and improves the final loss. All of the tensor products have learnable weights, meaning that every pair of input irreps that contribute to one output irrep has a learnable scalar parameter that is optimised during training. We use the AdamW 81 optimiser to minimise the loss, which is the mean squared error between predicted and training Hubbard parameters.
@@
To train and validate our ENN model, we curated a dataset comprising materials with diverse crystal structures. This dataset is constructed based on an investigation of various Li-ion cathode materials, covering olivine-type, 18 spinel-type, 52 and layered-type 82 structures at different Li concentrations, employing the self-consistent DFT+$U$ and DFT+$U$+$V$ protocols outlined above. Additionally, we include materials such as tunnel- and rutile-type MnO2 83, 51 and perovskite-type rare-earth nickelates 53 in our dataset. The crystal and electronic structure details, magnetic ordering, and various other properties of these materials are extensively documented in their respective publications. These are the compounds containing TM elements Ni, Mn, and Fe, and a list of all these materials is provided in table 2. While our current dataset is relatively small in terms of material count, ongoing efforts are directed towards a high-throughput exploration involving hundreds of materials, which will serve as an expanded training and validation set for a future, general-purpose, model. This initial dataset demonstrates the efficacy of our ML model, showcasing its ability to predict Hubbard parameters even with limited data (see Table 3). Despite the size of the dataset, the model exhibits accurate predictions, setting the stage for further refinement and expansion with the upcoming comprehensive investigation.
@@
Total 12 28
Table 2: A list of the materials that are used to train and validate the ML model. The last column shows the total number of materials. For the olivine-type materials $x=0,0.25,0.50,0.75,0$, while for the spinel- and layered-type materials $x=0,1$.
For each material in our dataset, SC Hubbard $U$ and $V$ parameters are computed using the self-consistent protocol illustrated in fig. 1. We include all $V$ parameters for Hubbard-active atom pairs whose DFPT values are greater than $0.3$ eV, in in constrast to Ref. 64 where only $V$ for the nearest neighbor couples were included in the training of their ML model. Throughout the iterative process, all intermediate converged occupation matrices and Hubbard parameters are systematically saved. The dataset for training $V$ comprises approximately seven times more data points than that for $U$ as $V$ is an atom-pair quantity, in contrast to the local and consequently sparser nature of $U$. For each material and chemical element, this data is randomly split, holding back 20% from each material for validation.
@@
Table 3: Chemical elements and the total number of data points for both on-site $U$ and inter-site $V$ Hubbard parameters determined using the SC protocol illustrated in fig. 1. These data points are aggregated across all the materials listed in table 2 that contain the respective chemical element. Numbers in brackets indicate the unique data points after de-duplication (see text).
In the first-principles calculations, certain input Hubbard parameters ($U_{\mathrm{in}}$ and $V_{\mathrm{in}}$) yield a specific ground state and occupation matrices, and the response of the system due to a perturbation through DFPT provides corresponding output Hubbard parameters ($U_{\mathrm{out}}$ and $V_{\mathrm{out}}$). We use these quantities to predict either the final self-consistent Hubbard parameter values, or the result of a single DFPT calculation to test various aspects of the model (see details for each experiment below). It is worth noting that many of the structures contain symmetry-equivalent sites, and correlations between results from successive self-consistent steps are prevalent. To address this issue, we implement a de-duplication procedure aimed at mitigating potential leakage of identical data from the training to the validation set. This involves calculating a symmetry-invariant distance between all pairs of ENN inputs for each atomic species in each material, which is then used to cluster duplicates based on a specified distance threshold (see Supplementary Section 4). In random train/validate splits, we sample from these clusters rather than individual data points, ensuring that only one example from each cluster is included in the validation set.
Figure 3: Parity plots showing the prediction accuracy on an unseen validation dataset. The energies in the legend are the RMSE categorized by element(s) and the overall RMSE across all elements. All attributes listed in table 1 are used as inputs for the ML model.
@@
First, we demonstrate the model’s ability to directly learn the final SC Hubbard parameters for the TM elements across all the materials in our dataset. This is achieved using all attributes listed in table 1 as inputs for our ML model. Figure 3 shows the parity plots obtained for the on-site $U$ and inter-site $V$ Hubbard parameters. The mean absolute relative error (MARE) over all species consistently remains below 3% and 5% for the $U$ and $V$ parameters, respectively (per-species distributions are reported in Sec. S5 in the SI). The result for Hubbard $U$ is particularly promising, considering the relatively small amount of available training data. The higher MARE obtained for $V$ is likely due to the significantly increased number of degrees of freedom in its model, owing to the involvement of pairs of atoms. It is worth noting that in order to attain this accuracy in predicting the $U$ and $V$ values, utilizing a relatively small batch size of $8-16$ was beneficial for preventing overfitting.
Second, we examine the extent to which our ML model depends upon the input Hubbard parameters for its prediction accuracy. To investigate this, we repeat the same numerical experiment as above, eliminating the input Hubbard parameters ($U_{\text{in}}$ and $V_{\text{in}}$) from our model. As detailed in Sec. S6 in the SI, the parity plots closely resemble those in fig. 3. The reduction in accuracy of our ML model resulting from this simplification is relatively small, namely the overall root mean square error (RMSE) is increases by 29 and 18 meV for the on-site $U$ and inter-site $V$, respectively. This finding suggests that the input Hubbard parameters may not be the most critical input attributes for our ML model, and that the input occupation matrices contain sufficient information to make low-error predictions.
@@
Table 4: Comparison of the open-circuit voltages $\Phi$ (in V) and magnetic moments for TM elements (in $\mu_{\mathrm{B}}$) for LixMnPO4 and LixFePO4 computed within DFT+$U$+$V$, with $U$ and $V$ obtained from first principles using DFPT and predicted using the ML model. The voltages are computed using the total energy differences for the Li concentrations $x=0$ and $x=1$, while the magnetic moments are computed as the trace of the difference between the spin-up and spin-down occupation matrices. 18 $\Delta\Phi$ and $\Delta m$ are the relative differences between the voltages and magnetic moments based on DFPT and ML, respectively.
In addition to evaluating the model accuracy, it is also important to evaluate its impact on downstream properties, such as voltages in Li-ion battery cathode materials. Table 4 shows a comparison of calculated open-circuit voltage (OCV) 18 $\Phi$ and magnetic moments within the DFT+$U$+$V$ framework, using first-principles SC Hubbard parameters from DFPT and those predicted by our ML model. Such OCV have been calculated using standard definitions 84, 85, which, when employed in the present context, imply that the energies of different Li concentrations are evaluated using different Hubbard parameters. This approach was used and discussed in more detail in previous works 86, 18, 83, while our goal here is to showcase the consistency between the results obtained with first-principles Hubbard parameters and the ones calculated through the ML model. We find that the differences between the computed and predicted $U$ and $V$ values are less than $0.12$ eV for LixMnPO4 and LixFePO4 ($x=0$ and $x=1$). Generally, such a small variation in the values of Hubbard parameters has a negligible impact on the vast majority of various physical and chemical properties of materials. The differences observed in the OCV and magnetic moments in Table 4 are indeed negligible, indicating the accuracy and reliability of the ML-predicted Hubbard parameters.
@@
In the preceding section, we assessed the ML model’s performance in predicting final self-consistent Hubbard parameters using training data that consists of a subset of DFPT calculations from all iterations of the SC protocol (see fig. 1). However, this approach relies on conducting numerous computationally intensive DFPT calculations to generate these training data (often $2-5$ but occasionally up to $10$ per material to reach self-consistency).
In this experiment, we instead investigate how well the model can predict the results of a single DFPT calculation as a function of how many self-consistent iterations are performed in generating training data. This task is more targeted towards high-throughput applications where data generation for model training could be more efficiently performed by sampling the first few DFPT calculations in the self-consistent workflows of many materials rather than the highly-correlated DFPT calculations of full workflows for a few materials. Effectively, this approach aims to stop the self-consistent cycle early and use a model trained on the calculations already performed to refine the not-yet-self-consistent result to be closer to self-consistency, essentially for free. We investigate this use case by training the model on only the first $N_{\mathrm{iter}}-1$ DFPT results, and asking it to predict the outcome of the next DFPT calculation and study how the model performs for different values of $N_{\mathrm{iter}}$.
Figure 4: Evaluation of model performance using a reduced number of self-consistent steps. RMSE for all materials as a function of the number of iterations $N_{\mathrm{iter}}$ in the SC protocol (see fig. 1). At each $N_{\mathrm{iter}}$ the plotted reference value is the RMSE between the DFPT result obtained from the previous SC iteration and the current one (i.e. a measure of how much the Hubbard $U$ changes by doing one more iteration). The ML value shows the prediction made from having trained on results from all the preceding $N_{\mathrm{iter}}-1$ iterations and predicting the $N_{\mathrm{iter}}$th value. This helps to answer the question of how much the ML model can improve the Hubbard parameter, at essentially no cost, when limiting the number of DFPT steps as might be done during the high-throughput screening.
@@
Figure 4 shows the average RMSE over all elements as the number of SC iterations used for training $N_{\mathrm{iter}}$ increases. The reference data correspond to RMSE evaluated using DFPT calculations, i.e. excluding the ML component. This RMSE is computed by determining the difference between the Hubbard parameters at the current $N_{\mathrm{iter}}$th and the previous $(N_{\mathrm{iter}}-1)$th iteration. It is evident that the reference RMSE decreases non-monotonically, which was observed in previous studies, 48 and eventually diminishes at large $N_{\mathrm{iter}}$. It is noteworthy that the optimal $N_{\mathrm{iter}}$ for achieving self-consistency in the SC protocol varies for different materials listed in table 2. Therefore, for larger $N_{\mathrm{iter}}$ values in fig. 4, fewer data points are available compared to smaller $N_{\mathrm{iter}}$ values.
@@
The ML-based RMSE in fig. 4 at the $N_{\mathrm{iter}}$th iteration is computed by training the model on data from all previous $N_{\mathrm{iter}}-1$ iterations and predicting the Hubbard parameters for the $N_{\mathrm{iter}}$th iteration. Section S7 in the SI contains parity plots for various values of $N_{\mathrm{iter}}$, ranging from 2 to 7, alongside the reference results broken down by element as well as the results for the training set. It is evident from fig. 4 that the ML-based RMSE steadily decreases, and for all $N_{\mathrm{iter}}$ values except $N_{\mathrm{iter}}=6$, it is smaller than the RMSE of the reference dataset. This result shows that our ML model can improve upon a DFPT-based SC result that was terminated early, particularly during the first few iterations, thereby facilitating faster convergence of Hubbard parameters. For instance, the reference value for $N_{\mathrm{iter}}=2$ indicates that on average, a single-shot DFPT calculation yields Hubbard parameters with an RMSE of $\sim 1$ eV, while the ML model trained on single-shot data can reduce this error to under $0.8$ eV. Incorporating additional $N_{\mathrm{iter}}$ iterations in the ML training data leads to steady improvements, although the relative performance compared to the reference gradually diminishes until reaching the floor of the model’s accuracy. This floor is partly determined by the convergence of the underlying DFPT-based Hubbard parameters in the SC protocol, conducted with a tolerance of approximately $\Delta=0.01-0.1$ eV (see fig. 1). Therefore, the most significant computational savings in predicting the SC Hubbard parameters can be achieved using our ML model, which acts as a surrogate, after only a few iterations in the DFPT-based SC protocol. This substantially reduces the overall computational cost and renders it a feasible approach, particularly for high-throughput screening scenarios where a priori training data is unavailable for all materials.
@@
Figure 5: Evaluation of model transferability to unseen electronic structures. RMSE as a function of $N_{\mathrm{c}}$ (the number of Li concentrations) for three olivines. The model is trained using $N_{\mathrm{c}}$ concentrations and validated on the remaining $5-N_{\mathrm{c}}$. Error bars indicate the range of RMSE values obtained by considering various permutations of concentrations of the training and validation datasets, large dots represent the average, small dots show individual results and experiments that only contain training data from a single oxidation state are labelled ($2+$ and $3+$). The reference is the RMSE computed by training the ML model on 80% of the Hubbard parameters from all five concentrations, with validation performed on the remaining 20% after de-duplication. The blue line represents the mean results over three runs with different random initializations, the confidence interval shows the standard deviation.
The outcomes of our numerical experiments are shown in fig. 5. As a reference, we compute the RMSE individually for each material by training the ML model on 80% of the data from all five concentrations and validating it on the remaining 20% of the data after de-duplication. This yields RMSE values of 126, 182, and 420 meV for the Fe, Mn, and mixed Fe-Mn olivines, respectively (see the horizontal lines in fig. 5). Subsequently, we investigate the RMSE values for scenarios in which the ML model is trained on fewer concentrations, denoted as $N_{\mathrm{c}}$, to assess the sensitivity of the ML model to the amount of the training data and its transferability for predicting Hubbard parameters at other concentrations. To accomplish this, we train the ML model on $N_{\mathrm{c}}$ concentrations and validate it on the remaining $5-N_{\mathrm{c}}$ concentrations, and then we average out over all possible permutations of the concentrations between the training and validation datasets. The average RMSE values are represented by larger dots, smaller dots show each individual result, while the error bars in fig. 5 indicate the maximum and minimum RMSE values resulting from the various permutations. When training on data from a single concentration (Nc = 1) we have labelled the results from $x=0$ and $1$ which correspond to 3+ and 2+ formal OSs respectively, in these cases the model is being tested on OSs it has not seen during training. Four out of six of these give a lower RMSE than experiments where the training data included a mixture of 2/3+ states, suggesting that the model is capable of making meaningful predictions outside of the OS seen during training. More generally, as $N_{\mathrm{c}}$ increases, the average RMSE for each material decreases, reaching values of approximately 106, 301, and 652 meV at $N_{\mathrm{c}}=4$ for the Fe, Mn, and mixed Fe-Mn olivines, respectively.
To test whether our model can transfer between different crystal structures we perform an additional experiment where we isolate data from the tunnel ($\alpha$) and rutile ($\beta$) phases of MnO2, training on one and testing predictions on the other. Table 5 shows that the model can transfer well between the two phases, achieving RMSEs comparable to those found when training on 80% of the data from both phases and predicting on the rest.
Trained on RMSE (meV)
80% of $\alpha$ $\beta$-MnO2 144
$\alpha$-MnO2 132
$\beta$-MnO2 157
Table 5: Hubbard $U$ results when training on $\alpha$-MnO2 data and predicting on $\beta$-MnO2 and vice versa as well as a reference trained on 80% of all of the deduplicated data and predicting on the remainder. In this case, the model shows good transferability to the unseen crystal structures.
Overall, these experiments demonstrate that, depending on the desired accuracy, it is not necessary to include training data from the exact target system that the model is being evaluated on, and even using examples of one or two similar structures may be sufficient to extrapolate to other compositions or crystal structures. Given the computational cost of SC calculations, this can lead to a significant speedup, particularly in cases requiring larger supercells and various configurations. These findings demonstrate the good transferability of our ML model, indicating its potential utility and reliability for predicting Hubbard parameters in materials not included in the model’s training dataset.